(NH4)3PO4, ammonium phosphate


Part D 

(NH4)3PO4, ammonium phosphate 


Part E 

C17H19NO3. morphine
Answer #1

part D

mass percent can be calculated as follows

let for a substance x

mass percent of x = (mass of x in compound / total mass of compound) *100

(NH4)3 PO4

molecular mass of compound = molar mass of NH4*3 + atomic mass of P+ atomic mass of O*4

molar mass = 18*3 + 31 + 16*4

molar mass = 149 gram/mole.

mass % of N = (42/149)*100 = 28.18%

mass % Of H = (12/149)*100 = 8.05%

mass % of O = (64/149)*100= 42.95%

mass % of P = (31/149)*100 = 20.80%

part E

molar mass of C17H19NO3

= 12*17 + 19*1 + 14 + 16*3

= 204 + 19+ 14 + 48

= 285 gram/mol

mass % of C = (204 / 285 )*100= 72.34%

mass % of H = (19/ 285) *100 = 6.66%

mass % of N = (14/285)*100 = 4.91 %

mass % of O = (48/285)*100= 16.84 %

hope your problem has been solved

thank you